# The trade-off between diagonal and off-diagonal elements in the eigenstate thermalization hypothesis

### v2

@@ Abstract
To bypass using local observables as intermediate quantities in proving the eigenstate thermalization hypothesis (ETH), we have introduced an observable-independent measure of distinguishability. In this paper, we establish the connection between this measure and several other ETH measures in a more natural way. We first demonstrate a universal trade-off relation between the diagonal and off-diagonal elements of the measure. We then extend this discussion to eigenstate typicality and the average observable. This trade-off relationship reveals that the exponential growth of off-diagonal elements directly suppresses their own values, as well as indirectly suppressing the diagonal elements. This provides a new perspective on the physical mechanisms underlying ETH. Finally, through numerical calculations on a one-dimensional Ising spin chain, we explore various trade-off relationships and examine strong and weak ETH.
@@ I Introduction
The equilibration and thermalization of an isolated quantum system are crucial for understanding how quantum statistical mechanics emerges from unitary quantum mechanics. The eigenstate thermalization hypothesis (ETH) [1, 2] posits that the expectation values of quantum observables in an energy eigenstate should approximately match the thermal expectation values, offering a potential mechanism for the thermalization of such systems. In the original proposal by Deutsch and Srednicki [3, 4, 5], a random perturbation is introduced to a closed quantum system, and the ETH is expected to hold if the perturbed system becomes chaotic. However, this randomness complicates the understanding of its physical origins and mathematical description.
@@
In many studies of the ETH, such as [8, 9, 6, 10, 11, 14, 12, 13, 7], one does not assume random energy perturbations or random Hamiltonians but instead aim to derive statistical properties solely from quantum properties. These studies focus on demonstrating the so-called ”strong (weak) ETH”. The strong (weak) ETH states that all (or almost all) energy eigenstates exhibit thermal properties with respect to local observables, and it postulates that the local distinguishability between an eigenstate and a universal density matrix decreases exponentially with system size. It is generally believed that strong ETH does not apply to integrable or many-body localized systems. On the other hand, the weak ETH emphasizes eigenstate typicality, suggesting that the average distinguishability between an eigenstate and the canonical ensemble decreases algebraically with system size, and it can be proven for generic translationally invariant systems, including integrable systems [9, 11, 12].
@@
In many mathematical proofs of the ETH, such as [15, 11, 12], they typically define a distinguishability measure based on an arbitrary observable and then demonstrate that the distinguishability between the eigenstate and the canonical ensemble approaches zero in the thermodynamic limit, regardless of the observable’s specific form. This approach motivates us to develop a distinguishability measure independent of observables. The trace norm used in [14] serves as such a measure, but it is challenging to transform and less convenient for further theoretical analysis. In our previous work [16], we explored ETH using Belavkin-Staszewski relative entropy and the variance $V(\rho_{B},O^{B_{1}}_{ij})$. However, we did not fully appreciate the significance of certain formulas or conduct numerical calculations on specific models, leading to conclusions similar to the weak ETH. In this paper, we further investigate the properties of $V(\rho_{B},O^{B_{1}}_{ij})$ and conduct numerical verification with specific models.
We begin in Section II with preliminary results about ETH, subsystem ETH, weak ETH and ETH with the average observable. We show how the distinguishability measure $V(\rho_{B},O^{B_{1}}_{ij})$ relates to commonly used measures and demonstrate how it can be extended to the average observable. In Section III, we prove the existence of a universal trade-off equation for the diagonal and off-diagonal elements of $V(\rho_{B},O^{B_{1}}_{ij})$. This trade off relationship can be extended to the discussion of eigenstate typicality and the average observable. In Section IV, we examine the trade-off relation and the strong (weak) ETH by numerically simulating an Ising spin chain with a transverse and longitudinal magnetic field. Finally, in Section V, we conclude the paper and discuss related issues.
@@ II Preliminaries
In this section, we review the basics of ETH and subsystem ETH, including weak ETH with eigenstate typicality as discussed in [6, 10, 11, 7], and eigenstate thermalization with macroscopic observables as explored in [11, 12].
@@ II.1 ETH and subsystem ETH
Consider an isolated or closed quantum system $B$ with Hamiltonian $H$. Suppose the system has size $N$ and $H$ has eigenvectors |$\ket{E_{i}},i=1,2,...,D_{N}$ with energy eigenvalues $E_{i}$, such that $H\ket{E_{i}}=E_{i}\ket{E_{i}}$. For a few-body observable $A^{B_{1}}$, the measure
$d^{(1)}(\Pi^{ij}_{B},\rho_{B};A^{B_{1}}):=\lvert\text{Tr}[A^{B_{1}}(\ket{E_{i}}\bra{E_{j}}-\rho_{B}\delta_{ij})]\rvert$ | (1)
quantifies the local (in)distinguishability between the energy eigenstates and the density matrix $\rho_{B}$. According to local ETH [1], this measure is expected to satisfy $d^{(1)}(\Pi^{ij}_{B},\rho_{B}(E);A^{B_{1}})\sim\mathcal{O}(e^{-S(E)/2})$, where $\rho_{B}(E)$ is some universal density matrix that changes slowly with energy $E=(E_{i}+E_{j})/2$.
@@
$d^{(2)}(\sigma^{B_{1}}_{ij},\rho_{B_{1}}):=\lVert\sigma^{B_{1}}_{ij}-\rho_{B_{1}}\delta_{ij}\rVert,$ (2)
@@
where $\sigma^{B_{1}}_{ij}=\text{Tr}_{\bar{B}_{1}}\ket{E_{i}}\bra{E_{j}}$ | is the reduced matrix on the subsystem $B_{1}$ and the trace distance $\lVert O\rVert=\frac{1}{2}\text{Tr}\sqrt{O^{\dagger}O}$. The measure (2) is used to study subsystem ETH, which posits that $d^{(2)}(\sigma^{B_{1}}_{ij},\rho_{B_{1}}(E))\sim\mathcal{O}(e^{-S(E)/2})$. Subsystem ETH, as defined by (2), is actually stronger than local ETH as described in (1). This is supported by the following inequality [14]
$d^{(1)}(\Pi^{i}_{B},\rho_{B};A_{B_{1}})\leqslant\sqrt{\lVert\sigma^{B_{1}}_{i}-\rho_{B_{1}}\rVert_{1}\text{Tr}[(\Pi^{i}_{B}+\rho_{B})A_{B_{1}}^{2}]},$ (3)
@@
where $\Pi^{i}_{B}=\ket{E_{i}}\bra{E_{i}}$ |, $\sigma_{i}=\sigma_{ii}$ and $\lVert\cdot\rVert_{1}=2\lVert\cdot\rVert$ is the Schatten $1$-norm (trace norm).
Although straightforward, the measure (2) is not in a very convenient form and is challenging to relate to other quantities. Using the Schatten $2$-norm instead of the trace norm yields a weaker condition, making it difficult to ensure that $d^{(1)}$ remains small. We introduce an alternative, less direct measure. By applying the rescaling map $\mathcal{J}^{\alpha}_{\rho}(\cdot):=\rho^{\alpha}(\cdot)\rho^{\alpha}$, we transform the elements in eq. 2 as follows:
$\sigma^{B_{1}}_{ij}-\rho_{B_{1}}\delta_{ij}=\mathcal{J}_{\rho_{B_{1}}}^{1/2}(O^{B_{1}}_{ij}-\braket{O^{B_{1}}_{ij}}),$ (4)
where $O^{B_{1}}_{ij}:=\mathcal{J}_{\rho_{B_{1}}}^{-1/2}(\sigma^{B_{1}}_{ij})$ and $\braket{O^{B_{1}}_{ij}}=\text{Tr}(O^{B_{1}}_{ij}\rho_{B})=\delta_{ij}$. Note that the rescaling map involves the inverse $\rho_{B_{1}}^{-1}$, requiring that the density matrix $\rho_{B_{1}}$ be strictly positive, which holds for the canonical and microcanonical ensembles considered here. Applying Hölder’s inequality, we obtain
$2d^{(2)}(\sigma^{B_{1}}_{ij},\rho_{B_{1}})=\lVert\mathcal{J}_{\rho_{B_{1}}}^{1/2}(O^{B_{1}}_{ij}-\braket{O^{B_{1}}_{ij}})\rVert_{1}$
$\leq\lVert(\rho_{B_{1}})^{1/2}(O^{B_{1}}_{ij}-\braket{O^{B_{1}}_{ij}})\rVert_{2}\lVert(\rho_{B_{1}})^{1/2}\rVert_{2}$
$=\sqrt{\text{Tr}[(O^{B_{1}}_{ij}-\braket{O^{B_{1}}_{ij}})^{\dagger}\rho_{B}(O^{B_{1}}_{ij}-\braket{O^{B_{1}}_{ij}})]}$
$=\sqrt{V(\rho_{B},O^{B_{1}}_{ij})}=:d^{(3)}(\sigma^{B_{1}}_{ij},\rho_{B_{1}}),$ (5)
where $\lVert T\rVert_{p}=\{\text{Tr}[(T^{\dagger}T)^{p/2}]\}^{1/p}$ denotes the Schatten $p$-norm and the variance is defined as:
$V(\rho,A):=\text{Tr}[(A-\braket{A})^{\dagger}\rho(A-\braket{A})].$ (6)
In this paper, we will focus on the measure $d^{(3)}$. As indicated by section II.1, $d^{(3)}$ serves as an upper bound for the other measures, making it the strongest ETH measure. Additionally, because variance is easily transformed, it facilitates connections with other quantities, which will be demonstrated later.
II.2 Weak ETH with eigenstate typicality
It is a strong assumption that all energy eigenstates satisfy the ETH. People are usually concerned with the Hilbert subspace within an ensemble. Consider the state of the total system $B$, denoted by $\rho_{B}=\sum_{j}p_{j}\Pi^{j}_{B}$. In many investigations of the “weak” ETH, the goal is to derive statistical properties and quantify probabilistic typicality or concentration with respect to the measure $d^{(1)}$. For this, the following expression is used [16]
$\Delta(\rho_{B},A)=\int[d^{(1)}(\Pi^{i}_{B},\rho_{B};A^{B_{1}})]^{2}p_{d},$ (7)

### v3

@@ Abstract
To bypass the reliance on local observables in verifying the eigenstate thermalization hypothesis (ETH), we introduce an observable-independent measure of distinguishability based on the variance of a rescaled local operator. We establish a universal trade-off relation between the diagonal and off-diagonal elements of this measure, rigorously connecting it to eigenstate typicality and spatially averaged observables. This trade-off reveals that exponential growth in the number of off-diagonal terms enforces their suppression, indirectly constraining diagonal deviations. Numerical simulations on a one-dimensional Ising spin chain with tunable transverse and longitudinal fields demonstrate stark contrasts between integrable and non-integrable regimes: While off-diagonal elements are universally suppressed with system size, diagonal suppression fails in integrable systems due to the absence of chaotic dynamics. Our results unify subsystem ETH, weak ETH, and macroscopic observables under a single framework, offering new insights into thermalization mechanisms.
@@ I Introduction
The equilibration and thermalization of isolated quantum systems are central to understanding how quantum statistical mechanics emerges from unitary quantum dynamics. The eigenstate thermalization hypothesis (ETH) [1, 2] proposes that expectation values of quantum observables in energy eigenstates should approximate thermal values, providing a potential thermalization mechanism. Deutsch’s foundational work [3] introduced random perturbations to closed quantum systems, suggesting ETH validity under chaotic dynamics. Srednicki’s later formulation using Berry’s conjecture [4, 5] also assumes that the eigenstates of the system (i.e., the wavefunctions) are superposition of random plane waves. While these randomness-based frameworks successfully derive thermalization, they risk obscuring the underlying physical origins of thermalization phenomena.
@@
Subsequent ETH studies [8, 9, 6, 10, 11, 14, 12, 13, 7] abandoned random Hamiltonian assumptions, instead deriving statistical properties from quantum principles. These investigations focus on demonstrating “strong (weak) ETH”. The strong (weak) ETH posits that all (or almost all) energy eigenstates exhibit thermal properties with respect to local observables, and it postulates that the local distinguishability between an eigenstate and a universal density matrix decreases exponentially with system size. It is generally believed that strong ETH does not apply to integrable or many-body localized systems. By contrast, the weak ETH emphasizes eigenstate typicality [6], where average distinguishability between an eigenstate and the canonical ensemble decays exponentially or algebraically with system size. Notably, weak ETH can be proven for generic translation-invariant systems, including integrable cases [9, 11, 12].
@@
Mathematical ETH proofs [15, 11, 12] typically define a distinguishability measure based on an arbitrary observable, and then demonstrate that the distinguishability between the eigenstate and the canonical ensemble approaches zero in the thermodynamic limit, regardless of the observable’s specific form. This motivates our development of an observable-independent measure. While trace norms [14] offer one such measure, their analytical complexity hinders deeper analysis. Our prior work [18] explored ETH using Belavkin-Staszewski relative entropy and the variance under a special local operator. However, we did not fully appreciate the significance of certain formulas, nor conduct numerical calculations on specific models, resulting in weak ETH conclusions with algebraic decay.
Here we significantly advance this approach by establishing fundamental properties of our variance measure and performing preliminary numerical verification. We begin in Section. II by bridging the variance measure to conventional ETH frameworks, demonstrating its superiority in unifying subsystem ETH, weak ETH, and macroscopic observables. In Section. III, we derive a universal trade-off equation that rigorously couples diagonal and off-diagonal elements, providing a model-agnostic foundation for analyzing ETH compliance. In Section IV, we numerically validate the trade-off relation using a tunable Ising spin chain, where varying the longitudinal field allows controlled switching between integrable and non-integrable regimes. By tracking the ratio between diagonal element and its corresponding average off-diagonal elements, we decode how integrability impedes diagonal suppression while leaving off-diagonal averages universally suppressed. Finally, in Section. V, we conclude the paper and discuss related issues.
@@ II Preliminaries
This section reviews fundamental concepts of ETH and subsystem ETH. We discuss weak ETH with eigenstate typicality as presented in [6, 10, 11, 7], along with eigenstate thermalization involving macroscopic observables from [11, 12].
@@ II.1 ETH and subsystem ETH
Consider a closed quantum system $B$ governed by Hamiltonian $H$, where $N$ denotes system size. Let {|$\{\ket{E_{i}}\}_{i=1}^{D_{N}}$ represent the Hamiltonian’s eigenstates with corresponding eigenvalues $E_{i}$, satisfying $H\ket{E_{i}}=E_{i}\ket{E_{i}}$. For a spatial subsystem $B_{1}$, one can define a local distinguishability measure for any few-body observable $A^{B_{1}}$ as
$|d^{(1)}(\Pi^{ij}_{B},\rho_{B};A^{B_{1}}):=\lvert\text{Tr}[A^{B_{1}}(\ket{E_{i}}\bra{E_{j}}-\rho_{B}\delta_{ij})]\rvert$ | (1)
to quantifies the local (in)distinguishability between the energy eigenstates and the density matrix $\rho_{B}$. Following local ETH [1], this measure follows
$d^{(1)}(\Pi^{ij}_{B},\rho_{B}(E);A^{B_{1}})=e^{-S(E)/2}\lvert f_{A}(E,\omega)R_{ij}\rvert,$ (2)
where $\rho_{B}(E)$ denotes a universal density matrix varying smoothly with mean energy $E:=(E_{i}+E_{j})/2$, $f_{A}(E,\omega)$ represents a smooth function of $E$ and energy difference $\omega:=E_{j}-E_{i}$, and $R_{ij}$ stands for a random variable with zero mean and unit variance.
@@
$d^{(2)}(\sigma_{B_{1}}^{ij},\rho_{B_{1}}):=\lVert\sigma_{B_{1}}^{ij}-\rho_{B_{1}}\delta_{ij}\rVert,$ (3)
@@
where $|\sigma_{B_{1}}^{ij}:=\text{Tr}_{\bar{B}_{1}}\Pi^{ij}_{B}=\text{Tr}_{\bar{B}_{1}}\ket{E_{i}}\bra{E_{j}}$ | denotes the reduced matrix on $B_{1}$, while $\rho_{B_{1}}:=\text{Tr}_{\bar{B}_{1}}\rho_{B}$ represents the subsystem’s reduced density matrix. The trace norm is defined as $\lVert O\rVert:=\frac{1}{2}\text{Tr}\sqrt{O^{\dagger}O}$. Subsystem ETH postulates $d^{(2)}(\sigma_{B_{1}}^{ij},\rho_{B_{1}}(E))\sim\mathcal{O}(e^{-S(E)/2})$, constituting a stronger condition than local ETH (2). This hierarchy is supported by the inequality [14]:
$d^{(1)}(\Pi^{i}_{B},\rho_{B};A_{B_{1}})\leqslant\sqrt{\lVert\sigma_{B_{1}}^{i}-\rho_{B_{1}}\rVert_{1}\text{Tr}[(\Pi^{i}_{B}+\rho_{B})A_{B_{1}}^{2}]},$ (4)
@@
where $|\Pi^{i}_{B}:=\ket{E_{i}}\bra{E_{i}}$ |, $\sigma^{i}:=\sigma^{ii}$, and Schatten $p$-norm $\lVert T\rVert_{p}:=\left\{\text{Tr}[(T^{\dagger}T)^{p/2}]\right\}^{1/p}$. The Schatten 111-norm (trace norm) relates to trace distance through $\lVert\cdot\rVert_{1}=2\lVert\cdot\rVert$.
While the measure (3) maintains conceptual simplicity, its practical implementation presents challenges due to analytical intractability and limited comparability with related quantities. The Schatten 2-norm (Frobenius norm) offers distinct advantages over the Schatten 1-norm (trace norm) through enhanced computational tractability stemming from its smoothness and differentiable properties. Substituting the trace distance in Eq. 3 with this alternative norm yields
$d^{\prime(2)}(\sigma_{B_{1}}^{ij},\rho_{B_{1}}):=\lVert\sigma_{B_{1}}^{ij}-\rho_{B_{1}}\delta_{ij}\rVert_{2}.$ (5)
While computationally favorable, this modified measure produces less stringent conditions and insufficiently constrains $d^{(1)}$ magnitude. We therefore propose an alternative approach employing a more sophisticated measure through rescaling transformations. Introducing the operator rescaling map $\mathcal{J}^{\alpha}_{\rho}(\cdot):=\rho^{\alpha}(\cdot)\rho^{\alpha}$, we reformulate the components of (3) as
$\sigma_{B_{1}}^{ij}-\rho_{B_{1}}\delta_{ij}=\mathcal{J}_{\rho_{B_{1}}}^{1/2}(O^{B_{1}}_{ij}-\braket{O^{B_{1}}_{ij}}_{\rho_{B}}),$ (6)
where the transformed local operator becomes $O^{B_{1}}_{ij}:=\mathcal{J}_{\rho_{B_{1}}}^{-1/2}(\sigma_{B_{1}}^{ij})$ with ensemble average $\braket{O^{B_{1}}_{ij}}_{\rho_{B}}:=\text{Tr}(O^{B_{1}}_{ij}\rho_{B})=\delta_{ij}$. This rescaling procedure necessitates strict positivity of $\rho_{B_{1}}$ due to inverse operator involvement – a condition naturally satisfied by canonical and microcanonical ensembles under consideration. Applying Hölder’s inequality establishes the crucial relationship
$2d^{(2)}(\sigma_{B_{1}}^{ij},\rho_{B_{1}})=\lVert\mathcal{J}_{\rho_{B_{1}}}^{1/2}(O^{B_{1}}_{ij}-\braket{O^{B_{1}}_{ij}}_{\rho_{B}})\rVert_{1}$
$\leq\lVert\rho_{B_{1}}^{1/2}(O^{B_{1}}_{ij}-\braket{O^{B_{1}}_{ij}}_{\rho_{B}})\rVert_{2}\times\lVert\rho_{B_{1}}^{1/2}\rVert_{2}$
$=\sqrt{\text{Tr}\left[(O^{B_{1}}_{ij}-\braket{O^{B_{1}}_{ij}}_{\rho_{B}})^{\dagger}\rho_{B}(O^{B_{1}}_{ij}-\braket{O^{B_{1}}_{ij}}_{\rho_{B}})\right]}$
$=\sqrt{V(\rho_{B},O^{B_{1}}_{ij})}=:d^{(3)}(\sigma_{B_{1}}^{ij},\rho_{B_{1}}),$ (7)
where the generalized variance measure is defined as:
$V(\rho,A):=\text{Tr}[(A-\braket{A}_{\rho})^{\dagger}\rho(A-\braket{A}_{\rho})]=\lVert\rho^{1/2}(A-\braket{A}_{\rho})\rVert^{2}_{2}.$ (8)
The variance expression in $d^{(3)}$ admits an operational reformulation through algebraic manipulation
$V(\rho_{B},O^{B_{1}}_{ij})=\lVert(\sigma^{ij}_{B_{1}}-\delta_{ij}\rho_{B_{1}})\rho^{-1/2}_{B_{1}}\rVert^{2}_{2}$
$=\text{Tr}(O^{B_{1}{\dagger}}_{ij}\rho_{B_{1}}O^{B_{1}}_{ij})-\delta_{ij}=\text{Tr}(\sigma^{ij}_{B_{1}}\rho^{-1}_{B_{1}}\sigma^{ij{\dagger}}_{B_{1}})-\delta_{ij}.$ (9)
A critical relationship emerges between measures $d^{(3)}$ and measure $d^{(1)}$, paralleling inequality (4):
$[d^{(1)}(\Pi^{ij}_{B},\rho_{B};A^{B_{1}})]^{2}=\lvert\text{Tr}[(\sigma^{ij}_{B_{1}}-\delta_{ij}\rho_{B_{1}})A^{B_{1}}]\rvert^{2}$
$=\lvert\text{Tr}[(\sigma^{ij}_{B_{1}}-\delta_{ij}\rho_{B_{1}})(A^{B_{1}}-\braket{A^{B_{1}}}_{\rho_{B}})]\rvert^{2}$
$\leq\lVert(\sigma^{ij}_{B_{1}}-\delta_{ij}\rho_{B_{1}})\rho^{-1/2}_{B_{1}}\rho^{1/2}_{B_{1}}(A^{B_{1}}-\braket{A^{B_{1}}}_{\rho_{B}})\rVert_{1}^{2}$
$\leq\lVert(\sigma^{ij}_{B_{1}}-\delta_{ij}\rho_{B_{1}})\rho^{-1/2}_{B_{1}}\rVert^{2}_{2}\times\lVert\rho^{1/2}_{B_{1}}(A^{B_{1}}-\braket{A^{B_{1}}}_{\rho_{B}})\rVert^{2}_{2}$
$=[d^{(3)}(\sigma_{B_{1}}^{ij},\rho_{B_{1}})]^{2}\times V(\rho_{B},A^{B_{1}}).$ (10)
In this inequality, the upper bound of $d^{(1)}$ is determined by two variances. One is state-dependent variance $V(\rho_{B},O^{B_{1}}_{ij})$, which is independent of the specific observable $A$ and determined by the intrinsic properties of the energy eigenstate. The other is the observable-dependent variance $V(\rho_{B},A^{B_{1}})$, which is independent of eigenstate and determined by the specific observable $A$. Although $d^{(3)}$ constitutes a stronger measure, the derived inequality Section II.1 can yield tighter bounds under certain circumstances (a particularly common scenario when considering the ETH). This can be demonstrated through Hölder’s inequality:
$\lVert(\sigma^{i}_{B_{1}}-\rho_{B_{1}})\rho^{-1/2}_{B_{1}}\rVert^{2}_{2}\leq\lVert\sigma^{i}_{B_{1}}-\rho_{B_{1}}\rVert^{2}_{1}\times\lVert\rho^{-1/2}_{B_{1}}\rVert^{2}_{\infty}.$ (11)
By comparing Eq. 4 with Section II.1, we observe that the upper bound provided by Section II.1 becomes tighter whenever $\lVert\sigma^{i}_{B_{1}}-\rho_{B_{1}}\rVert_{1}\times\lVert\rho^{-1/2}_{B_{1}}\rVert^{2}_{\infty}\leq 1$. This condition is naturally satisfied in ETH-considered systems: the trace norm $\lVert\sigma^{i}_{B_{1}}-\rho_{B_{1}}\rVert_{1}$ typically decays exponentially with system size, while the spectral norm $\lVert\rho^{-1/2}_{B_{1}}\rVert_{\infty}$ exhibits minimal scaling dependence. Numerical computations for specific models (as shown in Fig. 1) further validate this analysis. The plot demonstrates that for all energy eigenstates, the variance satisfies $V(\rho_{B},O^{B_{1}}_{ii})\sim\lVert\sigma^{i}_{B_{1}}-\rho_{B_{1}}\rVert^{2}_{1}<1$. This empirical evidence confirms that Section II.1 indeed provides a tighter bound compared to Eq. 4 in practical implementations.
The standard ETH framework in Eq. 2 imposes a critical constraint through the unit variance requirement of random variables $R_{ij}$, established via statistical averaging across an extensive eigenstate ensemble $\{i,j\}$. Combining Eq. 2 with Section II.1 yields the inequality
$\mathbb{E}\left[d^{(1)}(\Pi^{ij}_{B},\rho_{B}(E);A^{B_{1}})\right]^{2}=e^{-S(E)}\lvert f_{A}(E,\omega)\rvert^{2}$
$\leq\mathbb{E}\left[V(\rho_{B}(E),O^{B_{1}}_{ij})\right]\times V(\rho_{B}(E),A^{B_{1}}).$ (12)
According to this relationship, the variance characterization of $R_{ij}$ in the $d^{(1)}$ measure manifests as the ensemble average of $V(\rho_{B}(E),O^{B_{1}}_{ij})$ in the $d^{(3)}$ measure. Notably, the smooth functional behavior of $f_{A}(E,\omega)$ within the $d^{(1)}$ framework should be statistically recoverable from the $V(\rho_{B}(E),O^{B_{1}}_{ij})$ measurements. This correspondence implies that the characteristic features of the smooth function $f_{A}(E,\omega)$ can alternatively be extracted through systematic analysis of $V(\rho_{B}(E),O^{B_{1}}_{ij})$. Both the variance properties of $R_{ij}$ and the functional characteristics of $f_{A}(E,\omega)$ find significant applications in the investigation of two-point correlation functions [16, 17]. Research into these aspects presents a potential application of the properties of $V(\rho_{B}(E),O^{B_{1}}_{ij})$.
This work employs $d^{(3)}$ as the principal measure for ETH characterization, guided by its unique advantages: 1. Hierarchical Superiority: As rigorously established in Section II.1, $d^{(3)}$ provides a theoretical upper bound for preceding measures ($d^{(1)},d^{(2)},d^{\prime(2)}$), establishing it as the most stringent ETH criterion within this framework; 2. Operational Versatility: Operating within a mathematically tractable variance framework, $d^{(3)}$ enables systematic connections with thermodynamic observables through Section II.1; 3. Information Completeness: The variance structure $V(\rho_{B},O^{B_{1}}_{ij})$ intrinsically encodes both diagonal and off-diagonal correlations, providing unified access to ETH signatures across energy eigenstate ensembles. This tripartite rationale - bounding capability, analytical flexibility, and informational completeness - positions $d^{(3)}$ as our central investigative tool for subsequent thermalization analysis.
Table 1: Comparative Analysis of ETH Measurement Criteria
Property $d^{(1)}$ $d^{(2)}$ $d^{\prime(2)}$ $d^{(3)}$
Observable independence
$\rho_{B}$-dependent off-diagonals
Trade-off formulation Section III.1 N/A Section III.1 Eq. 22
Measurement hierarchy Weak Intermediate Weak Strong
State-independent trade-off
Legend: ✓ = Property satisfied, ✗ = Property not satisfied, N/A = Not applicable.
Measurement hierarchy reflects the relative strength in establishing ETH bounds, where stronger measures impose stricter constraints.
II.2 Effect of eigenbasis choice in degenerate systems
Consider $k$-fold degenerate eigenstates {|$\{\ket{\psi_{1}},\dots,\ket{\psi_{k}}\}$ with energy $E$. Under a unitary transformation $U$, any superposition state |$\ket{\psi^{\prime}}=U\ket{\psi}$ remains an eigenstate with the same energy. However, the measure $d(\sigma_{B_{1}}^{ij},\rho_{B_{1}})$ can change under the new eigenbasis {|$\{\ket{\psi^{\prime}}\}$. Here, we analyze how different choices of eigenbasis affect the measure $d^{(3)}$.
First, note that different eigenbases yield equivalent contributions to the reduced ensemble:
[…]
$|.\displaystyle=\text{Tr}_{\bar{B}_{1}}\left(\sum_{i=1}^{k}\ket{\psi^{\prime}_{i}}\bra{\psi^{\prime}_{i}}\right)=\sum_{i=1}^{k}\text{Tr}_{\bar{B}_{1}}\ket{\psi^{\prime}_{i}}\bra{\psi^{\prime}_{i}}.$ | . (13)
Consequently, the $\rho_{B_{1}}$ in measures $d^{(3)}$ remains invariant under superposition choices. As for the elements $|\sigma_{B_{1}}^{ij}=\text{Tr}_{\bar{B}_{1}}\ket{\psi_{i}}\bra{\psi_{j}}$ |, when considering different bases {|$\{\ket{\psi^{\prime}}\}$. The corresponding measure satisfies:
$V(\rho_{B},O^{\prime B_{1}}_{ij})$ $=\lVert\left(\sigma^{\prime ij}_{B_{1}}-\delta_{ij}\rho_{B_{1}}\right)\rho^{-1/2}_{B_{1}}\rVert^{2}_{2}$
$=\lVert\left(\sum_{ml}U_{im}U^{\dagger}_{lj}\sigma^{ml}_{B_{1}}-\delta_{ij}\rho_{B_{1}}\right)\rho^{-1/2}_{B_{1}}\rVert^{2}_{2}$
$\leq\delta_{ij}\sum_{m}|U_{im}|^{2}\lVert\left(\sigma^{mm}_{B_{1}}-\rho_{B_{1}}\right)\rho^{-1/2}_{B_{1}}\rVert^{2}_{2}$
$\quad+\sum_{m\neq l}|U_{im}U^{\dagger}_{lj}|\cdot\lVert\sigma^{ml}_{B_{1}}\rho^{-1/2}_{B_{1}}\rVert^{2}_{2}$
$\leq\delta_{ij}\max_{m}V(\rho_{B},O^{B_{1}}_{mm})+k\max_{m\neq l}V(\rho_{B},O^{B_{1}}_{ml}),$ (14)
where in the last inequality we have used the property of unitary transformations and
$\sum_{m\neq l}|U_{im}U^{\dagger}_{lj}|\leq(\sum_{m}|U_{im}|)(\sum_{l}|U_{jl}|)$
$\leq(\sum_{m}1)(\sum_{l}|U_{jl}|^{2})=k.$ (15)
As long as the degeneracy count $k$ grows significantly slower than the exponential decay constant of $d^{(3)}$, the superposition-induced discrepancies in the measures become negligible. Similar bounds apply to measures $d^{(2)}$, $d^{\prime(2)}$, and $d^{(1)}$, which can be analyzed using a method similar to Section II.2. Notably, $d^{(1)}$, being observable-dependent, may exhibit distinct behavior since degenerate states may remain indistinguishable for specific observables.
II.3 Weak ETH with eigenstate typicality
While theoretically robust, the strong ETH formulation may impose excessive constraints in practical scenarios. This limitation becomes particularly evident when examining the standard expression (2), which regulates only the variance of $R_{ij}$ through ensemble averaging, leaving maximal fluctuation magnitudes unbounded. Direct characterization of $R_{ij}$’s intrinsic stochastic properties presents significant analytical hurdles due to its implicit state dependence. A more operationally viable approach emerges through statistical ensemble analysis of observable concentration.
Consider an energy eigenstate ensemble for composite system $B$ described by density matrix $\rho_{B}=\sum_{j}p_{j}\Pi^{j}_{B}$. The “weak” ETH framework transitions focus to establishing statistical regularity and quantifying observable typicality through probabilistic measures. This probabilistic perspective employs the concentration measure [18]:
$\Delta(\rho_{B},A^{B_{1}}):=\int[d^{(1)}(\Pi^{i}_{B},\rho_{B};A^{B_{1}})]^{2}p_{d},$ (16)